There are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Sawz Logging Co. The company paid an average of $300 for each Douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy?
Answer by checkley75(3666)

850=D+P D=850-P NOW SUBSTITUTE (850-P) FOR D IN THE SECOND EQUATION
217,500=300D+225P
217,500=300(850-P)+225P
217,500=255,000-300P+225P
300P-225P=255,000-217,500
75P=37,500
P=37,500/75
P=500 NUMBER OF PONDEROSA PINES
850-500=350 NUMBER OF DOUGLAS FIRS
PROOF
217,500=300*350+225*500
217,500=105,000+112,500
217,500=217,500